hoodie:

If f(x) = 8 – 10x and g(x) = 5x + 4, what is the value of (fg)(–2)? –196 –168 22 78

Astrid1:

What we are going to do is replace and move the variables in this function. \[( f g ) ( − 2 ) = f ( − 2 ) × g ( − 2 ) .\] We can calculate each of these smaller bits using the definitions of the functions you gave in the question with x= - 2 \[f ( − 2 ) = 8 − 10 ( − 2 ) = 8 + 20 = 28\]\[g ( − 2 ) = 5 ( − 2 ) + 4 = − 10 + 4 = − 6\] Now, the equation becomes: \[( f g ) ( − 2 ) = 28 × − 6 = − 168\] -168 would be your answer.